OpenStudy (anonymous):

solve for the variable x- 1/8 = 3/4

OpenStudy (anonymous):

you could just convert the fractions to have the same demoninator of 8 so x-1/8=6/8 then add 1/8 to both sides x=7/8

OpenStudy (anonymous):

x=7/8